# Asymmetric Errors

### v1

@@ A.1 The Dimidiated Gaussian
$\sigma^{2}_{+}={\sigma_{x}^{+}}^{2}+{\sigma_{y}^{+}}^{2}\qquad\sigma_{\pm}^{x}={\sigma_{x}^{+}}^{2}+{\sigma_{y}^{-}}^{2}\qquad\sigma_{\mp}^{2}={\sigma_{x}^{-}}^{2}+{\sigma_{y}^{+}}^{2}\qquad\sigma_{-}^{2}={\sigma_{x}^{-}}^{2}+{\sigma_{y}^{-}}^{2}$ (23)
@@
$p_{z}(z)=\phi\left({z\over\sigma_{\mp}}\right)\Phi\left({-z\sigma_{x}^{-}\over\sigma_{y}^{+}\sigma_{\mp}}\right)+\phi\left({z\over\sigma_{+}}\right)\left[\Phi\left({z\sigma_{y}^{+}\over\sigma_{x}^{+}\sigma_{+}}\right)-\Phi\left({-z\sigma_{x}^{+}\over\sigma_{y}^{+}\sigma_{+}}\right)\right]+\phi\left({z\over\sigma_{\pm}}\right)\left[1-\Phi\left({z\sigma_{y}^{-}\over\sigma_{x}^{+}\sigma_{\pm}}\right)\right]$ (24)
@@
$p_{z}(z)=\hfill\phi({z\over\sigma_{\mp}})\Phi\left({z\sigma_{y}^{+}\over\sigma_{x}^{-}\sigma_{\mp}}\right)+\phi({z\over\sigma_{-}})\left[\Phi\left({-z\sigma_{x}^{-}\over\sigma_{y}^{-}\sigma_{-}}\right)-\Phi\left({z\sigma_{y}^{-}\over\sigma_{x}^{-}\sigma_{-}}\right)\right]+\phi\left({z\over\sigma_{\pm}}\right)\left[1-\Phi\left({-z\sigma_{x}^{+}\over\sigma_{y}^{-}\sigma_{\pm}}\right)\right]$ (25)
An example is shown in Figure 19. Two quantities $x$ and $y$, both with large positive skewness, ${\sigma^{+}}=1.5,{\sigma^{-}}=0.5$, are combined. The black curve shows the convolution of the two dimidiated Gaussians, using the above equations: if two dimidiated Gaussians are combined this is the ‘correct’ pdf. The red curve shows the curve from the ‘usual procedure’ of combining positive and negative sigmas separately and agreement is poor. The green curve shows the dimidiated Gaussian whose first 3 moments match that of the convolution, and gives much better agreement. Note that the central value (the median) has shifted, also that the asymmetry is reduced by the combination, in accordance with the Central Limit Theorem: the green curve has ${\sigma^{+}}=1.93,{\sigma^{-}}=0.97$.
@@ A.2 The Distorted Gaussian
Figure 20: The Distorted Gaussian. The left hand plot shows the results of a simple OPAT analysis, as before, and the proposed approximation. The right hand plot shows the pdf projected on the vertical axis that would come from a Gaussian on the horizontal axis.
The parameters $a$ and $b$ are given very simply from the quantile parameters: $a={1\over 2}({\sigma^{+}}+{\sigma^{-}})$ and $b={1\over 2}({\sigma^{+}}-{\sigma^{-}})$.
The pdf is again the probability transform of the unit Gaussian, but using a single quadratic transform rather than two piecewise linear ones. Although this is more plausible than the two straight lines of the dimidiated Gaussian, care needs to be taken to include both arms of the parabola in evaluating the pdf. The support for $p(R)$ is limited by the minimum (or, for a negative skewness, the maximum) of the parabola. Where this occurs there is a Jacobian peak, and this appears in figure around $R=22$.
If $\nu$ is distributed according to the unit Gaussian with $\mu=0$ and $\sigma=1$, $\phi(\nu)$, and $x(\nu)=x_{0}+a\nu+b\nu^{2}$, then the pdf for $x$ is $p(x)={{\phi}(\nu)\over|x^{\prime}(\nu)|}={{\phi}({\nu})\over|a+2b\nu|}$, where $a$ and $b$ are as given in Equation (26). As this is a quadratic, there are two values of $\nu$ for a given $x$: $\nu={\sqrt{a^{2}+4bx}-a\over 2b}$ and ${-\sqrt{a^{2}+4bx}-a\over 2b}$. Except for small asymmetries and/or small deviations, where the contribution from the second can be neglected, the pdf for $x$ is the sum of the two.
@@
From these, if desired, the other quantile parameters are just given by ${\sigma^{+}}=a+b,{\sigma^{-}}=a-b$.
@@ A.3 The “Railway” Gaussian
$R=\begin{cases}L(\nu,-1-h_{l})&\text{for }\nu\leq-\!1-h_{l}\\ T(\nu,-1,-h_{l})&\text{for }-\!1-h_{l}\leq\nu\leq-1\\ M+a\nu+b\nu^{2}&\text{for }-\!1\leq\nu\leq 1\\ T(\nu,1,h_{r})&\text{for }1\leq\nu\leq 1+h_{r}\\ L(\nu,1+h_{r})&\text{for }\nu\geq 1+h_{r}\end{cases}$ (29)
@@
According to Eq. 29, for the right transition region $[1,1+h_{r}]$ with $\nu_{0}=1$, we must set $f(\nu_{0})=M+1+b$, $f^{\prime}(\nu_{0})=a+2b$, $f^{\prime\prime}(\nu_{0})=2b$. The transition curve in the $[-\!1-h_{l},-1]$ left transition region is similarly parameterized by the function and its derivatives at $\nu_{0}=-1$. An example railway coordinate transformation and the corresponding density is shown in Fig. 21.
Figure 21: The panel on the left illustrates the coordinate transformation used to construct a railway Gaussian with parameters $M=0$, $\sigma^{+}=1$, $\sigma^{-}=0.6$, $h_{l}=h_{r}=1$. The panel on the right shows the corresponding density.
@@
The moments of the railway Gaussian have to be evaluated numerically. For numerical reasons, the values of $h$ used by the software are restricted to the range $[0.1,10.0]$.
Note that, in general, the railway coordinate transformation is allowed to be non-monotonous (for example, when $\sigma^{+}$ and $\sigma^{-}$ have different signs). In this case, similar to the distorted Gaussian, the railway Gaussian will have a semi-infinite support and will exhibit a Jacobian spike.
@@ A.4 The Double Cubic Gaussian
$R=M+\begin{cases}\frac{1}{4}(5{\sigma^{-}}-{\sigma^{+}})(\nu+1)-{\sigma^{-}}&\text{for }\nu\leq-\!1\\ \nu\left(\frac{1}{4}\left(\nu^{2}+3\nu+2\right)({\sigma^{+}}-{\sigma^{-}})+{\sigma^{-}}\right)&\text{for }-\!1\leq\nu\leq 0\\ \nu\left(\frac{1}{4}\left(\nu^{2}-3\nu+2\right)({\sigma^{-}}-{\sigma^{+}})+{\sigma^{+}}\right)&\text{for }0\leq\nu\leq 1\\ \frac{1}{4}(5{\sigma^{+}}-{\sigma^{-}})(\nu-1)+{\sigma^{+}}&\text{for }\nu\geq 1\end{cases}$ (31)
The dependence of $R$ on $\nu$ is continuous together with its first two derivatives, while the third derivative exhibits discontinuities at $\nu=-1,0,1$. An example double cubic coordinate transformation and the corresponding density are shown in Fig. 22.
Figure 22: The panel on the left illustrates the coordinate transformation used to construct a double cubic Gaussian with parameters $M=0$, $\sigma^{+}=1$, $\sigma^{-}=0.6$. The panel on the right shows the corresponding density.
@@ A.5 The Symmetric Beta Gaussian
$b(x,p,h)=\frac{1}{A}\frac{d^{2}R(\nu)}{d\nu^{2}}=\begin{cases}0&\text{for }|x|\geq h\\ \left(1-\left(\frac{x}{h}\right)^{2}\right)^{p}&\text{for }|x|<h\end{cases}$ (33)
is a shifted and scaled (as well as unnormalized) density of the symmetric beta distribution. $h$ is a positive width parameter. In order to maintain the continuity of the transformation second derivative and to enable efficient numerical calculation of the cumulants, the power parameter $p$ is expected to be a small positive integer (our software is restricted to $p\in\{1,...,20\}$). For given values of $p$, $h$, and $M$, parameters $A$ and $k$ are chosen to satisfy the conditions $R(-1)=M-{\sigma^{-}}$ and $R(1)=M+{\sigma^{+}}$, while the condition $R(0)=M$ is satisfied by Eq. 32 automatically.
The transformation defined by Eq. 32 is continuous together with its $p+1$ leading derivatives. This is reflected in the smoothness of the $R$ distribution whose density will possess at least $p$ continuous derivatives at all inner points of its support (the Jacobian spike, if present, lies at the support boundary). In the limit $h\rightarrow 0$ the distribution of $R$ tends to the dimidiated Gaussian, and in the limit $h\rightarrow\infty$ it becomes the distorted Gaussian. Practically usable values of $h$ are between 0.1 and 10.0 or so, values outside this range can lead to numerical instabilities in the software.
Examples of the transformations generated by Eq. 32, together with the corresponding densities, are illustrated in Fig. 23 for a number of different $p$ and $h$ settings.
Figure 23: The panel on the left illustrates the coordinate transformations used to construct the symmetric beta Gaussian densities shown on the right. In all cases, $M=0$, $\sigma^{+}=1$, $\sigma^{-}=0.6$.
@@ A.6 The Quantile Variable Width Gaussian
The quantile variable width (QVW) Gaussian distribution is a special case of the simple Q-normal distribution proposed in [ref:keelin2011]. It is based on the following idea. Suppose, $Q(y)$ is the quantile function (i.e., the inverse cumulative distribution function) of the standard normal. Then $\mu+\sigma\,Q(y)$ is the quantile function of the Gaussian distribution with mean $\mu$ and standard deviation $\sigma$. We can also define a quantile function which looks like $\mu+\sigma(y)\,Q(y)$. In some sense, distribution defined by this quantile function can be interpreted as a Gaussian with variable width, $\sigma(y)$. The width in this case depends on the cumulative probability rather than on the coordinate. The QVW Gaussian distribution is employing the following simple parameterization:
$\sigma(y)=\sigma_{0}\left[1+a\left(y-\frac{1}{2}\right)\right],$ (34)
where $a$ is a shape parameter (asymmetry). The complete quantile function is then
$q(y)=\mu_{0}+\sigma_{0}\left[1+a\left(y-\frac{1}{2}\right)\right]Q(y).$ (35)
The density that corresponds to the cdf value $y$ is, of course, $\left(\frac{dq(y)}{dy}\right)^{-1}$. The $y$ value corresponding to a given density argument $x$ has to be determined by solving the equation $q(y)=x$ numerically.
@@
Here, $\phi(x)$ is the standard normal density, and $\Phi(x)$ is the standard normal cdf. In particular, $M_{11}=\frac{1}{2\sqrt{\pi}}$, $M_{22}=\frac{\sqrt{3}+2\pi}{6\pi}$, $M_{31}=M_{32}=\frac{5}{4\sqrt{\pi}}$. Multiple precision numerical evaluation of $M_{33}$ gives 0.6751064260945980674284983.
An example QVW Gaussian density and the corresponding transformation of the standard normal are shown in Fig. 24
Figure 24: The panel on the right illustrates the QVW Gaussian density which corresponds to the asymmetric uncertainty result $0^{+1.0}_{-0.6}$. The panel on the left shows the corresponding coordinate transformation given by $x=q(\Phi(z))$.
@@ A.7 The Fechner distribution
The probability density function of the Fechner (a.k.a. split-normal) distribution is given by
$q(x)=\begin{cases}\sqrt{\frac{2}{\pi}}\frac{1}{\sigma_{1}+\sigma_{2}}\exp\left(-\frac{1}{2}\left(\frac{x-m}{\sigma_{1}}\right)^{2}\right)&\text{for }x\leq m,\\ \sqrt{\frac{2}{\pi}}\frac{1}{\sigma_{1}+\sigma_{2}}\exp\left(-\frac{1}{2}\left(\frac{x-m}{\sigma_{2}}\right)^{2}\right)&\text{for }x\geq m,\\ \end{cases}$ (38)
@@
$\begin{split}\mu&=m+\sqrt{\frac{2}{\pi}}(\sigma_{2}-\sigma_{1})\\ \sigma^{2}&=\left(1-\frac{2}{\pi}\right)(\sigma_{2}-\sigma_{1})^{2}+\sigma_{1}\sigma_{2}\\ \gamma&=\sqrt{\frac{2}{\pi}}(\sigma_{2}-\sigma_{1})\left[\left(\frac{4}{\pi}-1\right)(\sigma_{2}-\sigma_{1})^{2}+\sigma_{1}\sigma_{2}\right]\end{split}$ (39)
The largest possible normalized skewness, $s=\gamma/\sigma^{3/2}$, is reached for $\sigma_{1}=0$ and $\sigma_{2}>0$. It corresponds to the $s$ of the half-Gaussian, $\sqrt{2}(4-\pi)(\pi-2)^{-3/2}$. This limits the asymmetry that can be represented by the Fechner distribution, $A=\frac{\sigma^{+}-\sigma^{-}}{\sigma^{+}+\sigma^{-}}$, by about 0.21564027.
Construction of the Fechner distribution from moments or from quantiles has to be performed numerically. An example Fechner probability density function and the corresponding transformation of the standard normal are shown in Fig. 25.
Figure 25: The panel on the right illustrates the probability density function of the Fechner distribution which corresponds to the asymmetric uncertainty result $0^{+1.0}_{-0.7}$ (the result $0^{+1.0}_{-0.6}$ for which a number of other distributions is plotted in this Appendix can not be represented by Fechner). The panel on the left shows the corresponding coordinate transformation given by $x=q(\Phi(z))$.

### v2

@@ A.1.1 Convolution of two dimidiated Gaussians
$\sigma^{2}_{+}={\sigma_{x}^{+}}^{2}+{\sigma_{y}^{+}}^{2},\qquad\sigma_{\pm}^{x}={\sigma_{x}^{+}}^{2}+{\sigma_{y}^{-}}^{2},\qquad\sigma_{\mp}^{2}={\sigma_{x}^{-}}^{2}+{\sigma_{y}^{+}}^{2},\qquad\sigma_{-}^{2}={\sigma_{x}^{-}}^{2}+{\sigma_{y}^{-}}^{2},$ (21)
@@
$p_{z}(z)=\phi\left({z\over\sigma_{\mp}}\right)\Phi\left({-z\sigma_{x}^{-}\over\sigma_{y}^{+}\sigma_{\mp}}\right)+\phi\left({z\over\sigma_{+}}\right)\left[\Phi\left({z\sigma_{y}^{+}\over\sigma_{x}^{+}\sigma_{+}}\right)-\Phi\left({-z\sigma_{x}^{+}\over\sigma_{y}^{+}\sigma_{+}}\right)\right]+\phi\left({z\over\sigma_{\pm}}\right)\left[1-\Phi\left({z\sigma_{y}^{-}\over\sigma_{x}^{+}\sigma_{\pm}}\right)\right],$ (22)
@@
$p_{z}(z)=\hfill\phi({z\over\sigma_{\mp}})\Phi\left({z\sigma_{y}^{+}\over\sigma_{x}^{-}\sigma_{\mp}}\right)+\phi({z\over\sigma_{-}})\left[\Phi\left({-z\sigma_{x}^{-}\over\sigma_{y}^{-}\sigma_{-}}\right)-\Phi\left({z\sigma_{y}^{-}\over\sigma_{x}^{-}\sigma_{-}}\right)\right]+\phi\left({z\over\sigma_{\pm}}\right)\left[1-\Phi\left({-z\sigma_{x}^{+}\over\sigma_{y}^{-}\sigma_{\pm}}\right)\right].$ (23)
An example is shown in Figure 19. Two quantities $x$ and $y$, both with large positive skewness, ${\sigma^{+}}=1.5,{\sigma^{-}}=0.5$, are combined. The black curve shows the convolution of the two dimidiated Gaussians, using the above equations: if two dimidiated Gaussians are combined this is the ‘correct’ pdf. The red curve shows the curve from the ‘wrong method’ of combining positive and negative sigmas separately and agreement is poor. The green curve shows the dimidiated Gaussian whose first 3 moments match that of the convolution, and gives much better agreement. Note that the central value (the median) has shifted, also that the asymmetry is reduced by the combination, in accordance with the Central Limit Theorem: the green curve has ${\sigma^{+}}=1.93,{\sigma^{-}}=0.97$.
@@ A.2 The distorted Gaussian
Figure 20: The Distorted Gaussian. The left hand plot shows the results of a simple OPAT analysis that gave a result ${25.0}^{+2.0}_{-1.5}$,as before, and the proposed approximation. The right hand plot shows the pdf projected on the vertical axis that would come from a Gaussian on the horizontal axis.
The parameters $a$ and $b$ are given very simply from the quantile parameters: $a={1\over 2}({\sigma^{+}}+{\sigma^{-}})$ and $b={1\over 2}({\sigma^{+}}-{\sigma^{-}})$. The pdf is again the probability transform of the unit Gaussian, but using a single quadratic transform rather than two piecewise linear ones. Although this is more plausible than the two straight lines of the dimidiated Gaussian, care needs to be taken to include both arms of the parabola in evaluating the pdf. The support for $p(R)$ is limited by the minimum (or, for a negative skewness, the maximum) of the parabola. Where this occurs there is a Jacobian peak, and this appears in Figure 20 around $R=22$.
If $\nu$ is distributed according to the unit Gaussian with $\mu=0$ and $\sigma=1$ and $x(\nu)=x_{0}+a\nu+b\nu^{2}$, then the pdf for $x$ is $p(x)={{\phi}(\nu)\over|x^{\prime}(\nu)|}={{\phi}({\nu})\over|a+2b\nu|}$, where $a$ and $b$ are as given in Equation (24), and $\phi(\nu)$ is the unit Gaussian. As this is a quadratic, there are two values of $\nu$ for a given $x$: $\nu={\sqrt{a^{2}+4bx}-a\over 2b}$ and ${-\sqrt{a^{2}+4bx}-a\over 2b}$. Except for small asymmetries and/or small deviations, where the contribution from the second can be neglected, the pdf for $x$ is the sum of the two.
@@
From these, if desired, the other parameters are just given by ${\sigma^{+}}=a+b,{\sigma^{-}}=a-b$.
Because of the second arm of the parabola, the parameters $M,{\sigma^{+}},{\sigma^{-}}$ do not give the exact quantiles for $R$. Because of this ambiguity the distorted Gaussian is treated slightly differently in the software packages: the R implementation uses the parameters described here and the Python/C++ implementation uses the parameters giving the exact quantiles.
@@ A.3 The “railway” Gaussian
$R=f(\nu)=\begin{cases}L(\nu,-1-h_{l})&\text{for }\nu\leq-\!1-h_{l}\\ T(\nu,-1,-h_{l})&\text{for }-\!1-h_{l}\leq\nu\leq-1\\ M+a\nu+b\nu^{2}&\text{for }-\!1\leq\nu\leq 1\\ T(\nu,1,h_{r})&\text{for }1\leq\nu\leq 1+h_{r}\\ L(\nu,1+h_{r})&\text{for }\nu\geq 1+h_{r}\end{cases},$ (27)
@@
According to Equation 27, for the right transition region $[1,1+h_{r}]$ with $\nu_{0}=1$, we must set $f(\nu_{0})=M+1+b$, $f^{\prime}(\nu_{0})=a+2b$, $f^{\prime\prime}(\nu_{0})=2b$. The transition curve in the $[-\!1-h_{l},-1]$ left transition region is similarly parameterized by the function and its derivatives at $\nu_{0}=-1$. An example railway coordinate transformation and the corresponding density is shown in Figure 21.
Figure 21: The Railway Gaussian. The panel on the left illustrates the coordinate transformation used to construct a railway Gaussian with parameters $M=0$, $\sigma^{+}=1$, $\sigma^{-}=0.6$, $h_{l}=h_{r}=1$. The panel on the right shows the corresponding density.
@@
The R code (Appendix D) evaluates the moments of the railway Gaussian algebraically, whereas in the python/C++ code (Appendix C) the moments are evaluated numerically and, for numerical reasons, the values of $h$ used are restricted to the range $[0.1,10.0]$.
Note that, in general, the railway coordinate transformation is allowed to be non-monotonic(for example, when $\sigma^{+}$ and $\sigma^{-}$ have different signs). In this case, similar to the distorted Gaussian, the railway Gaussian will have a semi-infinite support and will exhibit a Jacobian spike.
@@ A.4 The double cubic Gaussian
$R=M+\begin{cases}\frac{1}{4}(5{\sigma^{-}}-{\sigma^{+}})(\nu+1)-{\sigma^{-}}&\text{for }\nu\leq-\!1\\ \nu\left[\frac{1}{4}\left(\nu^{2}+3\nu+2\right)({\sigma^{+}}-{\sigma^{-}})+{\sigma^{-}}\right]&\text{for }-\!1\leq\nu\leq 0\\ \nu\left[\frac{1}{4}\left(\nu^{2}-3\nu+2\right)({\sigma^{-}}-{\sigma^{+}})+{\sigma^{+}}\right]&\text{for }0\leq\nu\leq 1\\ \frac{1}{4}(5{\sigma^{+}}-{\sigma^{-}})(\nu-1)+{\sigma^{+}}&\text{for }\nu\geq 1\end{cases}.$ (29)
The dependence of $R$ on $\nu$ is continuous together with its first two derivatives, while the third derivative exhibits discontinuities at $\nu=-1,0,1$. An example double cubic coordinate transformation and the corresponding density are shown in Figure 22.
Figure 22: The Double Cubic Gaussian: the panel on the left illustrates the coordinate transformation used to construct a double cubic Gaussian with parameters $M=0$, $\sigma^{+}=1$, $\sigma^{-}=0.6$. The panel on the right shows the corresponding density.
@@ A.5 The symmetric beta Gaussian
$b(x,p,h)=\frac{1}{A}\frac{d^{2}R(\nu)}{d\nu^{2}}=\begin{cases}0&\text{for }|x|\geq h\\ \left[1-\left(\frac{x}{h}\right)^{2}\right]^{p}&\text{for }|x|<h\end{cases}$ (31)
is a shifted and scaled (as well as unnormalized) density of the symmetric beta distribution. $h$ is a positive width parameter. In order to maintain the continuity of the transformation second derivative and to enable efficient numerical calculation of the cumulants, the power parameter $p$ is expected to be a small positive integer (the C++/python software is restricted to $p\in\{1,...,20\}$; it is not yet implemented in the R software). For given values of $p$, $h$, and $M$, parameters $A$ and $k$ are chosen to satisfy the conditions $R(-1)=M-{\sigma^{-}}$ and $R(1)=M+{\sigma^{+}}$, while the condition $R(0)=M$ is satisfied by Equation 30 automatically.
The transformation defined by Equation 30 is continuous together with its $p+1$ leading derivatives. This is reflected in the smoothness of the $R$ distribution whose density will possess at least $p$ continuous derivatives at all inner points of its support. (The Jacobian spike, if present, lies at the support boundary.) In the limit $h\rightarrow 0$ the distribution of $R$ tends to the dimidiated Gaussian, and in the limit $h\rightarrow\infty$ it becomes the distorted Gaussian. Practically usable values of $h$ are between 0.1 and 10.0 or so; values outside this range can lead to numerical instabilities in the software.
Examples of the transformations generated by Equation 30, together with the corresponding densities, are illustrated in Figure 23 for a number of different $p$ and $h$ settings.
Figure 23: The Symmetric Beta Gaussian: the panel on the left illustrates the coordinate transformations used to construct densities shown on the right. In all cases, $M=0$, $\sigma^{+}=1$, $\sigma^{-}=0.6$.
@@ A.6 The quantile variable width Gaussian
The quantile variable width (QVW) Gaussian distribution is a special case of the simple Q-normal distribution proposed in [21]. It is based on the following idea. Suppose, $Q(z)$ is the quantile function (i.e., the inverse cumulative distribution function) of the standard normal. Then $\mu+\sigma\,Q(z)$ is the quantile function of the Gaussian distribution with mean $\mu$ and standard deviation $\sigma$. We can also define a quantile function which looks like $\mu+\sigma(z)\,Q(z)$. In a wider sense, the distribution defined by this quantile function can be interpreted as a Gaussian with variable width $\sigma(z)$. The width in this case depends on the cumulative probability rather than on the coordinate. The QVW Gaussian distribution is employing the following simple parameterization:
$\sigma(z)=\sigma_{0}\left[1+a\left(z-\frac{1}{2}\right)\right],$ (32)
where $a$ is a shape parameter governing the asymmetry. The complete quantile function is then
$q(z)=\mu_{0}+\sigma_{0}\left[1+a\left(z-\frac{1}{2}\right)\right]Q(z).$ (33)
The density that corresponds to the cdf value $z$ is, of course, $\left(\frac{dq(z)}{dz}\right)^{-1}$. The $z$ value corresponding to a given density argument $x$ has to be determined by solving the equation $q(z)=x$ numerically.
@@
Here, $\phi(z)$ is the standard normal density, and $\Phi(z)$ is the standard normal cdf. In particular, $M_{11}=\frac{1}{2\sqrt{\pi}}$, $M_{22}=\frac{\sqrt{3}+2\pi}{6\pi}$, $M_{31}=M_{32}=\frac{5}{4\sqrt{\pi}}$. Multiple precision numerical evaluation of $M_{33}$ gives 0.6751064260945980674284983.
An example QVW Gaussian density and the corresponding transformation of the standard normal are shown in Figure 24
Figure 24: The QVW Gaussian: the panel on the right illustrates the density which corresponds to the asymmetric uncertainty result $0^{+1.0}_{-0.6}$. The panel on the left shows the corresponding coordinate transformation given by $x=q(\Phi(z))$.
@@ A.7 The Fechner distribution
The probability density function of the Fechner (also known as the split-normal) distribution is given by
$P(x)=\begin{cases}\sqrt{\frac{2}{\pi}}\frac{1}{\sigma_{1}+\sigma_{2}}\exp\left(-\frac{1}{2}\left(\frac{x-m}{\sigma_{1}}\right)^{2}\right)&\text{for }x\leq m\\ \sqrt{\frac{2}{\pi}}\frac{1}{\sigma_{1}+\sigma_{2}}\exp\left(-\frac{1}{2}\left(\frac{x-m}{\sigma_{2}}\right)^{2}\right)&\text{for }x\geq m\\ \end{cases},$ (36)
@@
$\begin{split}\mu&=m+\sqrt{\frac{2}{\pi}}(\sigma_{2}-\sigma_{1})\\ \sigma^{2}&=\left(1-\frac{2}{\pi}\right)(\sigma_{2}-\sigma_{1})^{2}+\sigma_{1}\sigma_{2}\\ \gamma&=\sqrt{\frac{2}{\pi}}(\sigma_{2}-\sigma_{1})\left[\left(\frac{4}{\pi}-1\right)(\sigma_{2}-\sigma_{1})^{2}+\sigma_{1}\sigma_{2}\right]\end{split}.$ (37)
The largest possible normalized skewness, $\gamma/\sigma^{3/2}$, is reached for $\sigma_{1}=0$ and $\sigma_{2}>0$. It corresponds to the skewness of the half-Gaussian, $\sqrt{2}(4-\pi)(\pi-2)^{-3/2}$. This limits the asymmetry that can be represented by the Fechner distribution, $\frac{\sigma^{+}-\sigma^{-}}{\sigma^{+}+\sigma^{-}}$, to about 0.21564027.
Construction of the Fechner distribution from moments or from quantiles has to be performed numerically. An example Fechner probability density function and the corresponding transformation of the standard normal are shown in Figure 25.
Figure 25: The Fechner distribution: the panel on the right illustrates the probability density function which corresponds to the asymmetric uncertainty result $0^{+1.0}_{-0.7}$. (The result $0^{+1.0}_{-0.6}$ for which a number of other distributions is plotted in this appendix can not be represented by Fechner.) The panel on the left shows the corresponding coordinate transformation given by $x=q(\Phi(z))$.